# Palu: Compressing KV-Cache with Low-Rank Projection

### v1

@@ Appendix E Extended Experiments on LongBench and comparisons to quantization baselines
Table 5: Experiment results on Longbench. Average bit widths are estimated for each approach, assuming 10K context length. Each column represents the average score for the tasks of each type. Multi-QA and Single-QA denote Multi-Document and Single-Document Question Answering, respectively, while Summ refers to Summarization.
@@
Model Method Avg. Comp. Multi- Single- Summ Few-Shot Code Synthetic Avg.
Bits Ratio QA QA
@@
Table 6: Evaluation integrating LoRA with Palu on Llama2-7B.
@@
Table 7: Zero-shot accuracy of Palu on LLama2 models, with different decomposition strategies at 30% compression rate.
Model Method Zero-Shot Accuracy (%) $\uparrow$
OBQA Hella PIQA ARC-e ARC-c Wino Avg. Diff.
Llama2-7B Baseline 44.20 76.00 78.07 76.30 46.42 69.30 65.05 -
J-LRD 44.40 76.01 77.75 76.01 45.31 69.38 64.81 -0.24
G-LRD 45.20 76.07 78.56 75.93 44.97 68.75 64.91 -0.14
M-LRD 43.20 73.50 77.42 74.16 42.58 68.90 63.29 -1.76
Llama2-13B Baseline 45.20 79.39 79.11 79.42 49.06 72.38 67.43 -
J-LRD 45.20 79.58 80.57 79.67 49.40 71.74 67.69 +0.26
G-LRD 46.60 79.04 78.23 79.42 48.38 71.59 67.21 -0.22
M-LRD 45.40 78.57 78.78 78.83 47.61 72.06 66.88 -0.55
Appendix F Integrating Palu with LoRA Finetune
@@
Appendix G More Results on Zero-shot Accuracy
Following up Sec. 4.1, we further report the zero-shot evaluation results of Palu on the LLama2 family at 30% compression rate. As shown in Table. 7, we observe that Palu achieve nearly no accuracy drops using either J-LRD, G-LRD, or M-LRD. Thus, the users may adopt M-LRD first to optimize the efficiency further.
Table 8: Ablation study of low-rank decomposition group size on perplexity for the Llama2-7B model at a 50% compression rate using Wikitext-2.
@@ Appendix G More Results on Zero-shot Accuracy
Appendix H Ablation Study
H.1 Influence of Different Group Size
@@
H.2 Influence of Walsh-Hadamard Transform
@@
H.3 Automatic Rank Allocation vs. Uniform Rank Allocation
Table 10 presents the ablation study on the impact of different rank allocation schemes on the model’s accuracy. Applying rank searching results in a notable performance improvement. For instance, at a compression rate of 50%, there is a significant reduction in perplexity by 2.18. Fig. 7 visualizes the rank allocation across different transformer blocks for key and value projection layers. The results clearly demonstrate a non-uniform allocation result. Specifically, we observe that the value is generally allocated a higher rank compared to the key. Additionally, the first half of the layers are assigned higher ranks, indicating their greater importance in preserving model performance. This visualization underscores the effectiveness of our rank search algorithm in identifying and allocating appropriate ranks to different components, thereby optimizing the balance between compression and accuracy.
@@
Figure 7: Visualization of layer-wise low-rank compression rate on Llama2-7B with 50% of overall compression rate. Here, compression rates (i.e., rank) are allocated using the proposed Fisher Information-based automated rank allocation algorithm.
Table 10: Ablation study on w/ and w/o rank search. We use Llama2-7b and Wikitext-2 with sequence length 2048 as the benchmark.
@@
Appendix I Experiment Details
I.1 Zero-Shot Evaluation Details
@@
I.2 LongBench Evaluation Details
@@
During evaluation, we set the maximum sequence length to 8192 for the Mistral model and 15500 for Vicuna.

### v2

@@ 4.5 Latency Evaluation
Table 3: Experiment results on Longbench. Average bit widths are estimated for each approach, assuming 10K context length. Each column represents the average score for the tasks of each type.
@@
Model Method Avg. Comp. Multi- Single- Summa- Few-Shot Code Synthetic Avg.
Bits Ratio QA QA rization
@@
Figure 5: Normalized speedup for both the attention module and end-to-end model decoding. Solid lines represent exact measurements, while dashed lines indicate the FP16 baselines are out of memory, and the speedpus are compared the estimated baseline’s latency.
4.5.1 Speedups of Attention Module and End-to-end Decoding.
Attention module speedup.
We compare latency against standard attention without compression or quantization and show the speedups of Palu and KIVI-4-bit in Fig. 5 (a) and (b) for RoPE-based and non-RoPE attention. For RoPE-based attention, we applied our online reconstruction kernel for key and employed offline fusion for value as described in Sec 3.1.2.
As shown in Fig. 5 (a), Palu has minimal to no speedup when the sequence length is short, e.g. 4K. However, as sequence length increases, Palu delivers substantial performance gains. At 64K input length, Palu achieves a 1.89× speedup over the FP16 baseline when using low-rank projection alone.
By further applying 4-bit quantization to the Value states, the speedup rises to 2.91$\times$ for the same 64K context length, owing to our optimized low-precision kernel and reduced memory loading times. This performance notably surpasses KIVI-4-bit, which only achieves a 1.89× speedup at 64K, hindered by the overheads of its fined-grained group quantization. Notably, for RoPE-based attention, Palu-4-bit does not quantize key, as our online reconstruction kernel only supports FP16 precision for now.
For non-RoPE attention, we apply matrix fusion to both the Key and Value states (Eq. 6), effectively eliminating all reconstruction overhead. At a 64K sequence length with a 50% compression rate, Palu achieves a 2.20× speedup over the FP16 baseline. By further applying 4-bit quantization to both the Key and Value states, Palu boosts the speedup to 6.17× for 64K input length. These results demonstrate that combining low-rank compression and quantization significantly enhances inference efficiency, particularly in long-context scenarios.
End-to-end speedup.
We present the end-to-end speedups in Fig. 5 (c) and (d), measuring the latency of generating the next token at various input lengths and comparing the results to the FP16 baseline. Similar to the attention performance results, Palu shows minimal or no speedup for short sequences but delivers significant acceleration for longer sequences. Without quantization, Palu achieves up to 1.71× and 2.05× speedups for RoPE-based and non-RoPE models, respectively. With a 50% compression rate, Palu runs up to 32K input length on an RTX 4090 GPU. By incorporating 4-bit quantization, Palu handles even longer sequences and delivers 2.59× and 5.53× end-to-end speedups at a 64K sequence length. Palu integrated with quantization provides a substantial speed advantage over KIVI-4-bit, which only reaches 1.78× and 1.81× speedups at 32K sequence length for RoPE and non-RoPE scenarios, respectively, and is out-of-memory for longer sequences.
4.5.2 Kernel for RoPE-Based Attention Score
In this section, we evaluate the performance of our online reconstruction kernel for RoPE-based attention scores. We measure latency from the pre-RoPE query vector to post-GEMV attention score, and compare it with PyTorch’s GEMV, which is used in the baseline attention (see Fig. 2).
Figure 6: Speedup of Palu’s attention score kernel with online reconstruction.
We present speedups for group size 1, 4, and 32 at different sequence lengths in Fig. 6. For gs-32 (J-LRD), the highest accuracy decomposition, the high reconstruction cost causes a significant slowdown across all sequence lengths. For gs-1 (M-LRD), our kernel achieves up to a 3.56$\times$ speedup at sequence length 16K, showing strong performance when moderate accuracy loss is acceptable. For gs-4 (G-LRD), our kernel reaches up to 1.95× speedup. These results emphasize the need to explore various decomposition granularities for better accuracy and speed tradeoffs.
We also observe that speedup decreases for sequence lengths beyond 16K due to rising reconstruction costs, shifting the online reconstruction from memory- to compute-bound. A potential optimization is to quantize the decomposed weight matrices further and leverage high-throughput, low-precision hardware (e.g., INT4 Tensor Cores) for online reconstruction, which we leave for future work. Despite the speedup drop at longer lengths, Palu’s overall attention speedup increases with longer input, thanks to matrix fusion on the Value state and the reduced memory footprint.
5 Related Work
SVD for LLM Compression.
Several works have explored using SVD to compress LLMs. An early approach (Noach & Goldberg, 2020) applied standard SVD to weight matrices, resulting in significant compression errors. FWSVD (Hsu et al., 2022) addressed this by using Fisher information to prioritize parameters, while ASVD (Yuan et al., 2023) considered activation outliers. SVD-LLM (Wang et al., 2024) further minimized compression loss for each singular value. Unlike these methods, which compress model weights, Palu focuses on reducing KV-Cache size. A concurrent work, (Yu et al., 2024), also explores KV-Cache compression using low-rank projection, deriving low-rank matrices from KV-Cache with calibration data and grouping attention heads similarly to Palu’s G-LRD. However, it requires LoRA finetuning after decomposition. In contrast, Palu directly decomposes the weight matrix, preserving accuracy without finetuning, and introduces additional innovations, including rank search, quantization integration, and optimized GPU kernels.
KV-Cache Compression.
Quantization is a widely used technique for compressing KV-Cache. Atom (Zhao et al., 2023) applies simple per-token quantization, while WKVQuant (Yue et al., 2024) introduces a two-level scheme to enhance accuracy. KIVI (Liu et al., 2024b) uses per-channel and per-token quantization for Keys and Values, combined with fine-grained group quantization at group size 32. KVQuant (Hooper et al., 2024) employs a similar setup but incorporates non-uniform quantization and sparse matrices to handle outliers. On top of these approaches, GEAR adds a low-rank matrix to compensate for quantization errors. In Palu, we leverage low-rank techniques to exploit hidden dimension redundancy and, with simple per-token quantization, achieve outstanding compression results.
MLA.
The recently released DeepSeek-V2 model (DeepSeek-AI et al., 2024) introduces the MLA mechanism, which reduces KV-Cache size by down-projecting Key and Value to a low-rank space and reconstructing them to full rank at runtime. Although MLA may seem similar to Palu at a high level, particularly with J-LRD, our design and derivation processes are fundamentally different. Unlike MLA, a new attention mechanism requiring pre-training, Palu is specifically designed for post-training integration. Palu focuses on converting existing models with MHA or GQA to support low-rank compressed KV-Cache, preserving high accuracy while enhancing inference efficiency.
6 Conclusion
We introduce Palu, a novel KV-Cache compression framework that decomposes linear projection weight matrices and caches the compressed latent representations. We propose various optimizations, including group-head low-rank decomposition, automatic rank allocation algorithm, quantization compatibility enhancement, and customized kernels with operator fusion. With these optimizations, Palu can maintain accuracy while achieving significant memory reduction and high inference speedup. Experiments show that, with 50% low-rank compression and 4-bit quantization, Palu accelerates RoPE-based attention module by up to 2.91$\times$ and delivers up to 2.2$\times$ end-to-end speedup for the same RoPE-based model, while preserving strong accuracy on various benchmarks.
Appendix
Appendix A Quantization Basics
[…]
Appendix B Kernel Implementation Details
Figure 7: Illustration of our fused GPU kernel for computing attention scores with online reconstruction. In this figure, $\mathbf{q}$ represents the query vector, $\mathbf{H}$ denotes the low-rank compressed key states, and $\mathbf{B}$ stands for the reconstruction matrices.
Kernel for attention score calculation with reconstruction.
The central idea of Palu is to leverage low-rank latent representations to accelerate the attention mechanism by reducing data transfer overhead. Instead of working directly with the full-sized key matrix, we store and transfer a compressed low-rank latent representation, denoted as $\mathbf{H}\in\mathbb{R}^{L\times r}$. During computation, our custom GPU kernel performs an on-the-fly reconstruction using a reconstruction matrix $\mathbf{B}\in\mathbb{R}^{r\times d_{h}}$, producing a restored key matrix $\mathbf{K}\in\mathbb{R}^{L\times d_{h}}$, where $L$ is the sequence length, $d_{h}$ is the hidden dimension, and $r$ denote the remaining rank after performing low-rank projection. The query vector, represented as $\mathbf{q}\in\mathbb{R}^{1\times d_{h}}$, then multiplies with the reconstructed keys to obtain the attention scores.
To efficiently leverage parallelism, we perform tiling along the sequence length dimension $L$. Specifically, we split the sequence into smaller tiles of size $L_{\text{tile}}$, assigning each tile to a dedicated thread block. Each thread block independently reconstructs a submatrix $\mathbf{H}_{i}\in\mathbb{R}^{L_{\text{tile}}\times d_{h}}$ from the low-rank latent representation $\mathbf{H}$, then applies the positional embedding using RoPE, and finally performs the matrix-vector multiplication between $\mathbf{q}$ and $\mathbf{H}_{i}$ to produce partial attention scores. This design ensures that all intermediate computations, from reconstruction to embedding and final multiplication, remain entirely in on-chip memory (i.e., share memory), thus minimizing high-latency memory access and taking full advantage of the GPU’s parallel processing capabilities to achieve significant speedups.
Appendix C Discussion regarding memory usage
This issue primarily arises in J-LRD decomposition schemes, where the projections of all heads are decomposed jointly. In contrast, our M-LRD decomposition schemes and optimized G-LRD schemes involve non-square target matrices. For example, in the G-LRD scheme with a group size of 4, the target matrix is formed by concatenating the original projection matrices of each attention head in the group. In the Llama-2-7b model, with an embedding dimension of 4096 and head dimensions of 128, each projection matrix is 4096$\times$128, resulting in a concatenated matrix of size 4096$\times$512. In this case, the dimensions should be considered as $m=8n$. Applying the referenced equation $\frac{mr+nr}{mn}$ with $r=0.7n$, we find that $\frac{mr+nr}{mn}=0.7875$, indicating no additional storage cost and, in fact, achieving an additional 21.25% memory savings.
Appendix D Overall Memory Reduction
Figure 8: Total memory usage for various sequence lengths in Llama-2-7B. low-rank compression is at 50% for Palu.
Table 4: Evaluation integrating LoRA with Palu on Llama-2-7B.
@@ Appendix D Overall Memory Reduction
Appendix E Integrating Palu with LoRA Finetune
@@
Appendix F More Results on Zero-shot Accuracy
Following up Sec. 4.2, we further report the perplexity and zero-shot evaluation results of Palu on the Llama-2-13B at 50% compression rate. As shown in Table. 5, we observe that Palu achieve competitive accuracy drops around 3% or less across different using either J-LRD, G-LRD, or M-LRD. Thus, the users may adopt M-LRD first to optimize the efficiency further.
Table 5: Perplexity and zero-shot accuracy of Palu, with different decomposition strategies at 50%
Model Method Perplexity $\downarrow$ Zero-Shot Accuracy (%) $\uparrow$
Wiki2 C4 OBQA Hella PIQA ARC-e ARC-c Wino Avg.
Llama-2-13B Baseline 4.88 6.70 45.20 79.39 79.11 79.42 49.06 72.38 67.43
J-LRD 4.97 6.92 46.40 79.48 78.62 79.29 49.91 70.56 67.38
G-LRD 5.31 7.76 45.60 77.29 77.42 76.05 45.99 72.45 65.80
M-LRD 5.65 8.34 43.20 74.34 77.53 75.76 45.39 68.98 64.20
Table 6: Ablation study of low-rank decomposition group size on perplexity for the Llama-2-7B model at a 50% compression rate using Wikitext-2.
@@ Appendix F More Results on Zero-shot Accuracy
Appendix G Ablation Study
G.1 Influence of Different Group Size
@@
G.2 Influence of Walsh-Hadamard Transform
@@
G.3 Automatic Rank Allocation vs. Uniform Rank Allocation
Table 8 presents the ablation study on the impact of different rank allocation schemes on the model’s accuracy. Applying rank searching results in a notable performance improvement. For instance, at a compression rate of 50%, there is a significant reduction in perplexity by 2.18. Fig. 9 visualizes the rank allocation across different transformer blocks for key and value projection layers. The results clearly demonstrate a non-uniform allocation result. Specifically, we observe that the value is generally allocated a higher rank than the key. Additionally, the first half of the layers are assigned higher ranks, indicating their greater importance in preserving model performance. This visualization underscores the effectiveness of our rank search algorithm in identifying and allocating appropriate ranks to different components, thereby optimizing the balance between compression and accuracy.
@@
Figure 9: Visualization of layer-wise low-rank compression rate on Llama-2-7B with 50% of overall compression rate. Here, compression rates (i.e., rank) are allocated using the proposed Fisher Information-based automated rank allocation algorithm.
Table 8: Ablation study on w/ and w/o rank search. We use Llama-2-7b and Wikitext-2 with sequence length 2048 as the benchmark.
@@
Appendix H Experiment Details
H.1 Zero-Shot Evaluation Details
@@
H.2 LongBench Evaluation Details
@@
During the evaluation, we set the maximum sequence length to 31500 for both the Mistral and LongChat model.